Under identical conditions, separate samples of O2 and an unknown gas were allowed to effuse through identical membranes simultaneously. After a certain amount of time, it was found that 6.71 mL of O2 had passed through the membrane, but only 3.14 mL of of the unknown gas had passed through. What is the molar mass of the unknown gas? Number g/mol
Answer #1

From Graham’s law of effusion,

Rate is inversely proportional to square root of molar mass

Here:

rate(unknown)/rate(O2) = 3.14/6.71

rate(unknown)/rate(O2) = sqrt(M(O2)/M(unknown))

3.14/6.71 = sqrt(32/M(unknown))

0.468 = sqrt(32/M(unknown))

32/M(unknown) = 0.219

M(unknown) = 146 g/mol

Answer: 146 g/mol